b) A matter has a body-centered cubic unit cell with a length of 330 pm along an edge


and a density of 8.57 g cm^-3. Calculate the atomic mass of the matter.



c) Find the total energy of a moving object of mass 6.3 kg and momentum 11.5 x 10 kgms^-1

1
Expert's answer
2021-12-07T20:13:07-0500

b)


"(3.3\cdot10^{-10})^2=\frac{2M}{8.57\cdot6.023\cdot10^{23}}\\M=93"

c)


"E=\frac{p^2}{2m}\\E=\frac{(11.5\cdot10^6)^2}{2(6.3)}=1.05\cdot10^{13}\ J"